# GaussianFormer3D: Multi-Modal Gaussian-based Semantic Occupancy Prediction with 3D Deformable Attention

### v1

@@ 4.4 Ablation Study
Voxel-to-Gaussian Initialization. We first compare different levels of LiDAR features used for initializing Gaussian properties in Tab. 6(a). The improvement achieved with the multi-sweep voxel feature is significantly greater than that of the single-sweep point feature and the point cloud probabilistic modeling strategy used in GaussianFormer-2 [23], which validates the effectiveness of our proposed module. We further conduct an ablation study on the size of LiDAR voxel in initialization in Tab. 6(b). As the voxel size decreases, the model performance slightly improves. We choose $0.075\rm{m}\times 0.075\rm{m}\times 0.2\rm{m}$ as the final size.
@@
LiDAR-Guided 3D Deformable Attention. We first study the effect of the two-stage offset sampling strategy in Tab. 6(c). We observe that applying learnable offset sampling both before and after projection achieves higher performance than single-stage sampling, which validates that our two-stage sampling method can aggregate sufficient informative features for refining Gaussians. We also compare different feature aggregating methods for deformable attention in Tab. 6(d), including 3D deformable attention, 2D deformable attention with concatenated LiDAR sparse depth map and with completed dense depth map [29]. The results validate our design choice of 3D deformable attention.
4.5 Qualitative Results
@@
We visualize 3D Gaussians and occupancy to qualitatively verify the effectiveness of GaussianFormer3D for on-road scenes in Fig. 3. Our method can accurately predict both semantics and fine-grained geometry of the surrounding environments. In some cases, it even outperforms the ground truth by correctly completing occupancy in regions that lack semantic annotations. Qualitative results of our method on off-road scenes are given in Fig. 4. Our method is able to predict semantic occupancy for classes like mud and puddle, which are essential for achieving safe and effective off-road autonomous driving. We further compare our approach with GaussianFormer [25] in Fig. 5. The Gaussians in our method are more adaptive in scales and shapes, precisely appearing in the occupied regions of objects in both long-range and short-range areas, aided by the LiDAR sensor. Additionally, compared to voxel-based discretized approaches that train and predict at a fixed resolution, our method can predict multi-resolution semantic occupancy without additional training cost, attributed to the continuous property of Gaussians. This property enables more accurate and smoother prediction for certain areas when inferred at a higher resolution, as demonstrated in Fig. 6. Please see more qualitative results in the supplementary material.
5 Conclusion
@@
In this paper, we proposed GaussianFormer3D, a novel multi-modal semantic occupancy prediction framework that builds on 3D Gaussian scene representation and 3D deformable attention. We introduced a voxel-to-Gaussian initialization strategy to endow 3D Gaussians with accurate geometry priors from LiDAR data. We also designed a LiDAR-guided 3D deformable attention mechanism to refine 3D Gaussians with LiDAR-camera fusion feature in a lifted 3D space. Extensive experiments show the effectiveness of GaussianFormer3D in achieving accurate and fine-grained semantic occupancy prediction. In the future, we plan to explore the application of our multi-modal 3D Gaussian scene representation for multi-robot coordination.
Supplementary Material
Appendix A Related Work
Single-Modal Semantic Occupancy Prediction. Semantic occupancy prediction, also known as semantic scene completion, was firstly proposed to predict the occupancy and semantic states of all voxels within a pre-defined range for indoor scenarios [86, 34, 61, 9]. The vision-based and LiDAR-based pipelines are first explored, and further extended to outdoor environments, especially the autonomous driving scenes [4, 1, 39, 62]. The key to this challenging task is to learn an informative and compact scene representation, to which the common solutions mainly include 3D voxel-based representation [93, 77], 2D BEV-based representation [37, 79] and TPV-based (Tri-Perspective View) representation [22]. Voxel-based methods utilize discretized voxels to represent a 3D space, and transform semantic and geometric information from 2D images [5, 26, 36, 43, 48, 69, 70, 88, 18, 91, 47, 84, 53] or 3D point clouds [58, 76, 57, 10, 80, 73, 46, 6, 41, 75, 50] to voxel feature vectors. These approaches can effectively capture complicated structures but incur high storage and computational costs due to the dense nature of voxel-based representations. Moreover, a significant portion of these costs is unnecessary given the abundance of empty voxels in autonomous driving scenes. BEV-based representation has also been applied to semantic occupancy prediction [38, 82, 24] after its success in 3D object detection [37, 79]. TPV-based representation [22] extends BEV with two additional perpendicular planes and has been used in both vision-based [24, 22] and LiDAR-based [95] semantic occupancy networks. Although BEV-based and TPV-based methods mitigate the redundancy issue caused by the empty grids, their compression schemes inevitably lead to information loss. 3D Gaussian splatting [28] provides an expressive and compact scene representation, which has been explored in vision-based 3D semantic occupancy prediction [25, 23, 13]. GaussianFormer [25] represents an autonomous driving scene as a set of 3D Gaussians, designed to only model the object-occupied regions, which significantly improved prediction efficiency and reduced memory consumption.
Appendix B WildOcc Dataset
Figure 7: Distribution of categories across the WildOcc [83] train, validation, and test sets. The bar charts illustrate the percentage of each class within the respective dataset splits.
Figure 8: Qualitative results of GaussianFormer3D on the WildOcc [83] validation and test sets. In the off-road scenes, our model can still capture the geometry structures and semantic information of large continuous surfaces (rubble, grass) or small irregular objects (barrier, mud), as shown in the red boxes. Best viewed on screen and in color.
[…]
Based on the split files from RELLIS-3D [27], we filter out samples in the WildOcc that lack corresponding semantic occupancy labels, LiDAR poses, or ground truth depth, obtaining 7399, 1249 and 1399 samples for training, validation and testing, respectively. The test set originally contains 11 classes in total. However, to ensure comparability with the baselines in [83], we follow the same class selection strategy, considering semantic classes including grass, tree, bush, barrier, puddle, mud, and rubble, while grouping the remaining labels into the others category. The percentage distribution of each class is presented in Fig. 7.
A key challenge of this dataset is the significant distribution shift between the training/validation sets and the test set. Notably, as shown in Fig. 7, the barrier and rubble classes exhibit extremely low occurrence in the test set compared to other categories. Specifically, the test set contains only 273 and 233 voxels for barrier and rubble, respectively, accounting for merely 0.0008% and 0.0007% of the occupied voxels. This severe class imbalance poses challenges for model performance, as limited training data can hinder the accurate prediction of these underrepresented categories. Additionally, the insufficiency of captured features further complicates the prediction process, making it even more challenging to reliably infer these classes.
Appendix C Evaluation Metrics
For evaluation metrics, we utilize Intersection-over-Union (IoU) and mean Intersection-over-Union (mIoU) following MonoScene [5]. The IoU and mIoU can be computed as:
$\mathrm{IoU}$ $=\frac{TP_{\neq c_{0}}}{TP_{\neq c_{0}}+FP_{\neq c_{0}}+FN_{\neq c_{0}}},$ (7)
$\mathrm{mIoU}$ $=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{i\in\mathcal{C}^{\prime}}{\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}},$ (8)
where $TP$, $FP$, $FN$, $c_{0}$ and $\mathcal{C}^{\prime}$ represent the number of true positive, false positive, false negative predictions, the empty class and the set of non-empty classes respectively.
Figure 9: Qualitative results of GaussianFormer3D on the Occ3D [64] validation set. Our method is able to recover the fine-grained geometry structures and accurate semantics of the on-road scenes, and even outperforms the ground truth by completing the occupancy at some areas without annotations. Best viewed on screen and in color.
Figure 10: Multi-resolution occupancy predicted by the same 3D Gaussians on the SurroundOcc [70] validation set. The regions within the red boxes show that the high-resolution prediction generated by the same 3D Gaussians can achieve more accurate and smoother results, compared to the low-resolution one and the occupancy ground truth. Best viewed on screen and in color.
Appendix D Implementation Details
For multi-view images, we only apply photometric distortion augmentation, image normalization and image padding. During evaluation, we do not apply any test time augmentation technique. For the 3D deformable attention, DFA3D [31] has proved that the trilinear interpolation in the 3D feature space can be transformed into a depth-weighted bilinear interpolation, which means that in implementation, 3D deformable attention can be transformed into a depth-weighted 2D deformable attention to simultaneously maintain theoretical equivalence and improve efficiency. Hence we take advantage of the depth-weighted 2D deformable attention operator implemented in CUDA by DFA3D [31] to conduct 3D deformable attention.
Appendix E Supplementary Experiments
E.1 Qualitative Results
Qualitative results on the WildOcc [83] validation and test sets. We present visualization results of our model on the WildOcc [83] validation and test sets in Fig. 8. Two representative samples from both the validation and test sets are selected for illustration. Notably, all sequences in the validation and test sets are completely unseen during training. The class distribution has been reported in Appendix B. In the first validation sample (top row), our method successfully predicts the rubble region (highlighted in red boxes), which corresponds to the pile of rocks in the camera view. In the second validation sample, our model effectively captures the barrier structures in the scene. These results indicate that our model can learn semantic occupancy even with limited training data. In the test set samples (bottom two rows), our method accurately reconstructs large continuous surfaces, such as grass and trees, and effectively detects irregular objects and terrains, such as bush and mud, as shown in the red boxes. Overall, our predictions exhibit a high degree of similarity to the ground truth, especially in preserving the geometric structures of various terrain types, including grass, tree, rubble, and barrier.
Qualitative results on Occ3D [64] validation set. We provide visualization results of our model on the Occ3D [64] validation set in Fig. 9. Our method achieves fine-grained semantic occupancy prediction, and even occasionally outperforms the ground truth. Even though the occupancy ground truth is sparse and missing at some regions, our model is able to accurately recovering semantic information of these unlabeled areas.
Figure 11: Visualization comparison of different Gaussian initialization strategies on the SurroundOcc [70] validation set. With LiDAR voxel features as priors, our 3D Gaussians precisely appear at the regions of interests, which paves the way for the following 3D deformable attention-based Gaussian update.
[…]
Method Modality Sunny Rainy Day Night Sunny Rainy Day Night
GaussianFormer [25] C 29.6 27.5 30.3 19.5 18.9 18.0 19.2 9.3
Table 7: 3D semantic occupancy prediction results on SurroundOcc [70] validation set for different weather and lighting conditions. We quantitatively show the improvement made by GaussianFormer3D over the baseline GaussianFormer.
Multi-resolution semantic occupancy prediction on SurroundOcc validation set. We provide more qualitative results of multi-resolution semantic occupancy prediction in Fig. 10. As shown in red boxes in Fig. 10, we observe that the high-resolution prediction achieves more accurate and smoother occupancy at some regions. Although Gaussians are supervised with the fixed-resolution occupancy ground truth, multi-resolution predictions can still be generated by the same Gaussian representation due to its continuous modeling property. This advantage eradicates the need of training several models to predict occupancy of different resolutions, and keeps the high prediction accuracy with the same Gaussians, which significantly shortens the algorithm deployment time and saves the computation resources.
Visualization comparison of different Gaussian initialization strategies. We provide qualitative results of visualization comparison of different Gaussian initialization strategies in Fig. 11. Compared to random initialization, our designed voxel-to-Gaussian (V2G) initialization strategy is able to endow the 3D Gaussians with accurate position and geometry priors from the LiDAR data at the beginning of training, which paves the way for the following 3D deformable attention-based Gaussian refinement.
E.2 Quantitative Results
Model performance under different weather conditions. To break down the performance improvement under different weather conditions, we group the scenes in the nuScenes [4] dataset based on climate and lighting conditions. We demonstrate the results of both GaussianFormer3D and GaussianFormer [25] in Tab. 7. Compared to the baseline GaussianFormer, our approach presents increased performance across all weather conditions, which validates the effectiveness and robustness of our model. Due to the introduction of LiDAR sensor, GaussianFormer3D shows a significant performance improvement over the camera-only baseline under extreme climate (rainy) and low lighting condition (night). Both experiments are conducted with 25,600 Gaussians for a fair comparison.

### v2

@@ IV-D Ablation Study
Voxel-to-Gaussian Initialization. We first compare different levels of LiDAR features used for initializing Gaussian properties in Tab. VII(a). The improvement achieved with the multi-sweep voxel feature is significantly greater than that of the single-sweep point feature and the point cloud probabilistic modeling strategy used in GaussianFormer-2 [9], which validates the effectiveness of our proposed module. We further conduct an ablation study on the size of LiDAR voxel in initialization in Tab. VII(b). As the voxel size decreases, the model performance slightly improves. We choose $0.075\rm{m}\times 0.075\rm{m}\times 0.2\rm{m}$ as the final size.
@@
LiDAR-Guided 3D Deformable Attention. We first study the effect of the two-stage offset sampling strategy in Tab. VII(c). We observe that applying learnable offset sampling both before and after projection achieves higher performance than single-stage sampling, which validates our two-stage sampling method can aggregate sufficient informative features for refining Gaussians. We also compare different feature aggregating methods in Tab. VII(d), including 3D deformable attention, 2D deformable attention with concatenated LiDAR sparse depth map and with completed dense depth map [14]. The results validate our final choice.
IV-E Qualitative Results
@@
We visualize 3D Gaussians and occupancy to qualitatively verify the effectiveness of our method for on-road scenes in Fig. 4. Our method can accurately predict both semantics and fine-grained geometry of the surrounding environments. In some cases, it even outperforms the GT by correctly completing occupancy in regions that lack semantic annotations. Qualitative results of our method on off-road scenes are given in Fig. 5. Our method is able to predict semantic occupancy for classes like mud and puddle, which are essential for achieving safe and effective off-road autonomous driving. We further compare our approach with GaussianFormer [11] in Fig. 6. The Gaussians in our method are more adaptive in scales and shapes, precisely appearing in the occupied regions of objects in both long-range and short-range areas, aided by the LiDAR sensor. Additionally, compared to voxel-based discretized approaches that train and predict at a fixed resolution, our method can predict multi-resolution semantic occupancy without additional training cost, attributed to the continuous property of Gaussians. This property enables more accurate and smoother prediction for certain areas when inferred at a higher resolution, as demonstrated in Fig. 7.
V Conclusion
@@
In this paper, we proposed GaussianFormer3D, a novel multi-modal semantic occupancy prediction framework that builds on 3D Gaussian scene representation. We introduced a voxel-to-Gaussian initialization strategy to endow 3D Gaussians with accurate geometry priors from LiDAR data. We also designed a LiDAR-guided 3D deformable attention mechanism to refine 3D Gaussians with LiDAR-camera fusion feature. Extensive experiments show its effectiveness in achieving accurate and fine-grained semantic occupancy prediction. However, our model is limited to fully-supervised manner, which requires densely annotated occupancy labels for training. In the future, we will explore its self-supervised variant and its application for multi-robot coordination.